OpenStudy (anonymous):

solve 3x + 5y = 3 x + 2y = 0

OpenStudy (anonymous):

someone help me!

OpenStudy (anonymous):

3x + 5y = 3 --> 2 x + 2y = 0 x = -2y --> 1 sub 1 into 2 3(-2y) + 5y = 3 -6y + 5y = 3 -y = 3 y = -3 when y = -3 , x = -2(-3) = 6